# Zihao Zhang, Guodong Li, and Sihuang Hu

### v1

@@ 4 $(k+2,k,2)$ MDS array codes with optimal repair
Recall that an $(n=k+2,k,\ell=2)$ MDS array code comprises $n$ nodes, denoted by $C_{1},C_{2},\dots,C_{n}$, where each node is a column vector of length $\ell=2$. An MDS array code can be defined by the parity check sub-matrices $\mathbf{H}_{i}$, $i\in[n]$, and the parity-check equations (1) as described in Section 2. The defined code is MDS if and only if any two parity check sub-matrices $\mathbf{H}_{i},\mathbf{H}_{j}$ (where $1\leq i<j\leq n$) can form an invertible matrix. The linear repair process for each node $C_{i}$ can be described by a $2\times 4$ repair matrix $\mathbf{M}_{i}$ (where $i\in[n]$). The repair bandwidth and repair I/O for each node $C_{i}$ can be calculated using the formulas (3) and (4), respectively.
@@ 4.1 $(k+2,k,2)$ MDS array codes with optimal bandwidth
We give the parity check sub-matrix $\mathbf{H}_{i}$ and repair matrix $\mathbf{M}_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) in Table 1. The following lemma shows that the resulting code $\mathcal{C}_{1}$ is MDS.
$i$ $\mathbf{M}_{i}$ $\mathbf{H}_{i}$
@@
Table 1: Repair matrix $\mathbf{M}_{i}$ and parity check sub-matrix $\mathbf{H}_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) of code $\mathcal{C}_{1}$.
$\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ $j\in\mathcal{G}_{1}$ $j\in\mathcal{G}_{2}$ $j\in\mathcal{G}_{3}$ $j\in\mathcal{G}_{4}$
@@
Table 2: The value of $\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ for each $i,j\in[n]$ of code $\mathcal{C}_{1}$.
@@
We only need to show that each square matrix $\mathbf{H}_{i},\mathbf{H}_{j}]$ is invertible for all distinct indices $i,j$ (where $1\leq i<j\leq n$).
One key observation for the structure of the parity check sub-matrices in Table 1 is that any square matrix $\mathbf{H}_{i},\mathbf{H}_{j}]$ can be transformed into a block triangular matrix by an invertible transformation, and the blocks on the main block diagonal are invertible, which implies that the square matrix $\mathbf{H}_{i},\mathbf{H}_{j}]$ is also invertible.
If $i\in\mathcal{G}_{1}$ and $j\in\mathcal{G}_{2}$, we can add the second block row of $\mathbf{H}_{i},\mathbf{H}_{j}]$ to the first block row, and then permute the columns of the matrix to obtain a block triangular matrix. The procedure is as follows:
@@
Since the square matrix on the main block diagonal is invertible, the original matrix $\mathbf{H}_{i},\mathbf{H}_{j}]$ is also invertible. For other cases, we can directly permute the columns of the matrix to obtain a block triangular matrix. ∎
Repair bandwidth. Recall the linear repair process of each $C_{i}$ driven by the repair matrix $\mathbf{M}_{i}$ (where $i\in[n]$) described in Section 2. We can calculate the repair bandwidth for $C_{i}$ by the formulas (3).
Firstly, one can directly check from Table 1 that for each $i\in[n]$, $\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{i}\right)=2$. Thus, $\mathbf{M}_{i}$ can define a linear repair process for node $C_{i}$. Then we summarize the values of $\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ for all $i,j\in[n]$ in Table 2. Now we can calculate the repair bandwidth for each node $C_{i}$ (where $i\in[n]$) by the formula (3):
$\operatorname{BW}(\mathbf{M}_{i})$ $=\sum_{j=1}^{n}\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)-2$
@@ 4.2 $(k+2,k,2)$ MDS array codes with optimal repair I/O
We give the parity check sub-matrix $\mathbf{H}_{i}$ and repair matrix $\mathbf{M}_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) in Table 3. Lemma 19 shows that the resulting code $\mathcal{C}_{2}$ is MDS. We omit the proof of this lemma since it is similar to the proof of Lemma 18.
$i$ $\mathbf{M}_{i}$ $\mathbf{H}_{i}$
@@
Table 3: Repair matrix $\mathbf{M}_{i}$ and parity check sub-matrix $\mathbf{H}_{i}$ for each node $\mathcal{C}_{i}$ (where $1\leq i\leq n$) of code $\mathcal{C}_{2}$.
$\textnormal{nz}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ $j\in\mathcal{G}_{1}$ $j\in\mathcal{G}_{2}$ $j\in\mathcal{G}_{3}$
@@
Table 4: The value of $\textnormal{nz}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ for each $i,j\in[n]$ of code $\mathcal{C}_{2}$.
@@
Repair I/O. Recall the linear repair process of each $C_{i}$ driven by the repair matrix $\mathbf{M}_{i}$ (where $i\in[n]$) described in Section 2. We can calculate the repair I/O for $C_{i}$ using the formula (4).
Firstly, from Table  3, one can directly check that for each $i\in[n]$, $\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{i}\right)=2$. Thus, $\mathbf{M}_{i}$ can define a linear repair process for node $C_{i}$. Then we summarize the values of $\textnormal{nz}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)$ for all $i,j\in[n]$ in Table 4. Now we can calculate the repair I/O for each node $C_{i}$ (where $i\in[n]$) using the formula (4):
$\operatorname{IO}(\mathbf{M}_{i})$ $=\sum_{j=1}^{n}\textnormal{nz}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)-2$
@@ 5 Conclusion
We investigate lower bounds on the repair bandwidth and repair I/O for single-node failures in MDS array codes that have fixed sub-packetization. Using a linear-algebraic framework, we prove these bounds for codes with redundancy nk=2n-k=2 and sub-packetization $\ell=2$. We further present two explicit code constructions that meet the bounds, thereby establishing their tightness.

### v2

@@ 4 $(k+2,k,2)$ MDS Array Codes with Optimal Pepair
Recall that an $(n=k+2,k,\ell=2)$ MDS array code comprises $n$ nodes, denoted by $C_{1},C_{2},\dots,C_{n}$, where each node is a column vector of length $\ell=2$. An MDS array code can be defined by the parity check sub-matrices $H_{i}$, $i\in[n]$, and the parity-check equations (1) as described in Section 2. The defined code is MDS if and only if any two parity check sub-matrices $H_{i},H_{j}$ (where $1\leq i<j\leq n$) can form an invertible matrix. The linear repair process for each node $C_{i}$ can be described by a $2\times 4$ repair matrix $M_{i}$ (where $i\in[n]$). The repair bandwidth and repair I/O for each node $C_{i}$ can be calculated using the formulas (3) and (4), respectively.
@@ 4.1 $(k+2,k,2)$ MDS Array Codes with Optimal Bandwidth
We give the parity check sub-matrix $H_{i}$ and repair matrix $M_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) in Table 1. The following lemma shows that the resulting code $\mathcal{C}_{1}$ is MDS.
$i$ $M_{i}$ $H_{i}$
@@
Table 1: Repair matrix $M_{i}$ and parity check sub-matrix $H_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) of code $\mathcal{C}_{1}$.
$\textnormal{rank}\left(M_{i}H_{j}\right)$ $j\in\mathcal{G}_{1}$ $j\in\mathcal{G}_{2}$ $j\in\mathcal{G}_{3}$ $j\in\mathcal{G}_{4}$
@@
Table 2: The value of $\textnormal{rank}\left(M_{i}H_{j}\right)$ for each $i,j\in[n]$ of code $\mathcal{C}_{1}$.
@@
We only need to show that each square matrix $H_{i},H_{j}]$ is invertible for all distinct indices $i,j$ (where $1\leq i<j\leq n$).
One key observation for the structure of the parity check sub-matrices in Table 1 is that any square matrix $H_{i},H_{j}]$ can be transformed into a block triangular matrix by an invertible transformation, and the blocks on the main block diagonal are invertible, which implies that the square matrix $H_{i},H_{j}]$ is also invertible.
If $i\in\mathcal{G}_{1}$ and $j\in\mathcal{G}_{2}$, we can add the second block row of $H_{i},H_{j}]$ to the first block row, and then permute the columns of the matrix to obtain a block triangular matrix. The procedure is as follows:
@@
Since the square matrix on the main block diagonal is invertible, the original matrix $H_{i},H_{j}]$ is also invertible. For other cases, we can directly permute the columns of the matrix to obtain a block triangular matrix. ∎
Repair bandwidth. Recall the linear repair process of each $C_{i}$ driven by the repair matrix $M_{i}$ (where $i\in[n]$) described in Section 2. We can calculate the repair bandwidth for $C_{i}$ by the formulas (3).
Firstly, one can directly check from Table 1 that for each $i\in[n]$, $\textnormal{rank}\left(M_{i}H_{i}\right)=2$. Thus, $M_{i}$ can define a linear repair process for node $C_{i}$. Then we summarize the values of $\textnormal{rank}\left(M_{i}H_{j}\right)$ for all $i,j\in[n]$ in Table 2. Now we can calculate the repair bandwidth for each node $C_{i}$ (where $i\in[n]$) by the formula (3):
$\operatorname{BW}(M_{i})$ $=\sum_{j=1}^{n}\textnormal{rank}\left(M_{i}H_{j}\right)-2$
@@ 4.2 $(k+2,k,2)$ MDS Array Codes with Optimal Repair I/O
We give the parity check sub-matrix $H_{i}$ and repair matrix $M_{i}$ for each node $C_{i}$ (where $1\leq i\leq n$) in Table 3. Lemma 19 shows that the resulting code $\mathcal{C}_{2}$ is MDS. We omit the proof of this lemma since it is similar to the proof of Lemma 18.
$i$ $M_{i}$ $H_{i}$
@@
Table 3: Repair matrix $M_{i}$ and parity check sub-matrix $H_{i}$ for each node $\mathcal{C}_{i}$ (where $1\leq i\leq n$) of code $\mathcal{C}_{2}$.
$\textnormal{nz}\left(M_{i}H_{j}\right)$ $j\in\mathcal{G}_{1}$ $j\in\mathcal{G}_{2}$ $j\in\mathcal{G}_{3}$
@@
Table 4: The value of $\textnormal{nz}\left(M_{i}H_{j}\right)$ for each $i,j\in[n]$ of code $\mathcal{C}_{2}$.
@@
Repair I/O. Recall the linear repair process of each $C_{i}$ driven by the repair matrix $M_{i}$ (where $i\in[n]$) described in Section 2. We can calculate the repair I/O for $C_{i}$ using the formula (4).
Firstly, from Table 3, one can directly check that for each $i\in[n]$, $\textnormal{rank}\left(M_{i}H_{i}\right)=2$. Thus, $M_{i}$ can define a linear repair process for node $C_{i}$. Then we summarize the values of $\textnormal{nz}\left(M_{i}H_{j}\right)$ for all $i,j\in[n]$ in Table 4. Now we can calculate the repair I/O for each node $C_{i}$ (where $i\in[n]$) using the formula (4):
$\operatorname{IO}(M_{i})$ $=\sum_{j=1}^{n}\textnormal{nz}\left(M_{i}H_{j}\right)-2$
@@ 5 Conclusion
We investigate lower bounds on the repair bandwidth and repair I/O for single-node failures in MDS array codes with $r=2$ and $\ell=2$. Under these parameters, we map repair schemes to point sets on the projective line $\mathbb{P}^{1}$. Leveraging the sharply 3-transitive action of $\text{PGL}_{2}(\mathbb{F}_{q})$, we derive a lower bound on the repair bandwidth and provide explicit MDS array code constructions that meet this bound. Using a similar approach, we also establish a lower bound for repair I/O and provide an optimal construction.